Circular No. 7898

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002ct IN ANONYMOUS GALAXY M. Schwartz, Cottage Grove, OR; and W. Li, University of California

at Berkeley, report the LOTOSS (cf. IAUC 7514) discovery of an apparent supernova (mag ∼ 20.1) on unfiltered images taken with the 0.81-m Tenagra II telescope on May 10.3 and 12.2 UT. The new object is located at α = 12^{h}47^{m}22^{s}.53, δ = +28^{o}00′13′′.9 (equinox 2000.0), which is ∼ 13′′ west and 8′′ south of the nucleus of the apparent host galaxy. A Tenagra II image taken on Apr. 14.3 showed nothing at this position (limiting mag ∼ 22).

SUPERNOVA 2002cu NEAR NGC 6575 Li reports the LOTOSS discovery of a possible supernova on unfiltered

CCD images taken with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) on May 11.5 (mag ∼ 17.6) and 12.5 UT (mag ∼ 17.3). Located at α = 18^{h}10^{m}50^{s}.40, δ = +31^{o}06′32′′.3 (equinox 2000.0), SN 2002cu is 91′′.3 west and 26′′.0 south of the nucleus of NGC 6575. A KAIT image taken on May 5.5 showed nothing at this position (limiting mag ∼ 18.5).

T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2002cu, obtained by M. Calkins on May 13.45 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova near maximum. The spectrum appears very similar to that of SN 1999ac at maximum. The Si II (rest 635.5 nm) line is fairly broad (FWHM 11400 km/s) and would indicate an exceptional expansion velocity of 17400 km/s if the host galaxy is NGC 6575 (adopting the NASA/IPAC Extragalactic Database recession velocity of 6992 km/s).

COMETS C/1999 F3 AND C/1999 W2 (SOHO) Further to IAUC 7897, D. Hammer reports his measurements for ad-

ditional comets found by R. Kracht on SOHO archival C2 images (C/1999 F3 was also visible in C3 coronagraph images). The reduced astrometry and orbital elements by B. G. Marsden appear on the MPECs cited below. Comet C/1999 W2 is a member of the Kreutz group; C/1999 F3 belongs to the Meyer group (cf. IAUC 7832).

Comet  1999 UT  α_{2000}  δ_{2000}  MPEC
C/1999 F3 Mar. 17.238  23^{h}50.^{m}8  + 0^{o}08′  2002-H26
C/1999 W2 Nov. 26.439  16 07.0  −22 34  2002-G45
2002 May 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green